OpenStudy (anonymous):

Find the derivative; y=x^3+3^x

OpenStudy (anonymous):

\[\huge y=x^3+3^x\]

OpenStudy (blurbendy):

what are you trying to solve?

OpenStudy (anonymous):

Opps yes derivative

OpenStudy (anonymous):

Really? x^3 --> 3x^2

OpenStudy (anonymous):

\[3x^2\] for the first one

OpenStudy (anonymous):

then \[\frac{d}{dx}b^x=b^x\ln(b)\] finishes

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):